# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@ 1.2. The pinning model and weak coupling limit
(1.19) ${\mathbbm{E}}[\zeta_{n}]=0,\quad\sigma_{N}^{2}={\rm Var}(\zeta_{n}):=e^{\lambda(2\beta_{N})-2\lambda(\beta_{N})}-1.$
@@
(1.20) $\sigma_{N}^{2}=e^{\lambda(2\beta_{N})-2\lambda(\beta_{N})}-1=\frac{1}{R_{N}}\Big{(}1+\frac{\vartheta+o(1)}{\log N}\Big{)},\quad\text{for some fixed }\vartheta\in{\mathbbm{R}}.$
Using $\lambda(\beta)=\sum_{i=2}^{4}\frac{\kappa_{i}}{i!}\beta^{i}+O(\beta^{5})$ as $\beta\to 0$, where $\kappa_{i}$ is the $i$-th cumulant of $\omega$, a direct computation (see [CSZ19a, Appendix A.4] and [CSZ23, equation (3.12)]) yields that
(1.21) $\beta_{N}^{2}=R_{N}^{-1}+C_{1}(\omega)R_{N}^{-3/2}+C_{2}(\omega,\vartheta)R_{N}^{-2}+o(R_{N}^{-2}),$
@@
It is straightforward to see that CDPM $\mathscr{L}^{\vartheta}$ is translation invariant in law,
(1.22) $\mathscr{L}^{\vartheta}({\mathrm{d}}(s+{\rm a}),{\mathrm{d}}(t+{\rm a}))\overset{{\rm dist}}{=}\mathscr{L}^{\vartheta}({\mathrm{d}}s,{\mathrm{d}}t),\quad\forall\hskip 1.0pt{\rm a}\in\mathbb{R}.$
@@
(1.23) $\mathscr{L}^{\vartheta}({\mathrm{d}}({\rm a}s),{\mathrm{d}}({\rm a}t))\overset{{\rm dist}}{=}{\rm a}\hskip 1.0pt\mathscr{L}^{\vartheta+\log{\rm a}}({\mathrm{d}}s,{\mathrm{d}}t),\quad\forall\hskip 1.0pt{\rm a}>0.$
The above translation and scaling relations are analogous to those satisfied by the critical 2d stochastic heat flow [CSZ23, Theorem 9.2], and their proofs are exactly the same, which we leave to the readers.
@@
While the first and second moments assumptions on $S$ in (1.10) are virtual, we conjecture that the assumptions of vanishing third moment and finite fourth moment are technical, which are employed to establish a more refined bound, such as (4.25)-(4.27), and could be eliminated with a more careful analysis.
A general renewal process of the form (1.4) can be generated by Bessel random walks, which are a family of one-dimensional nearest-neighbor random walks with inhomogeneous increments (see [Ale11] for the choice of spatially dependent transition probabilities). For the critical case $\alpha=\frac{1}{2}$ with non-constant slowly varying function $L(\cdot)$, the transition probabilities are slightly inhomogeneous
@@
For two random walks $S$ and $\widetilde{S}$ both satisfying (1.10), it is not hard to see that the difference of their expected number of overlap $R_{N}-\widetilde{R}_{N}$ has a finite limit as $N$ tends to $\infty$. Denote the limit ${\rm a}({S,\widetilde{S}}):=\lim_{N\to\infty}(R_{N}-\widetilde{R}_{N})$. Suppose $\mathscr{L}^{\vartheta}({\mathrm{d}}s,{\mathrm{d}}t)$ is the limit w.r.t. $S$ and a choice of $\beta_{N}$ satisfying (1.20). Then, for $\widetilde{S}$ and the same $\beta_{N}$,
(1.24) $\widetilde{\sigma}_{N}^{2}=\frac{1}{\widetilde{R}_{N}}\Big{(}1+\frac{\widetilde{R}_{N}-R_{N}}{\log N/(2\pi)}\hskip 1.0pt\frac{\log N/(2\pi)}{R_{N}}\Big{)}\Big{(}1+\frac{\vartheta+o(1)}{\log N}\Big{)}=\frac{1}{\widetilde{R}_{N}}\Big{(}1+\frac{\vartheta-2\pi{\rm a}({S,\widetilde{S}})+o(1)}{\log N}\Big{)},$
@@
To show the convergence of ${\mathcal{Z}}_{N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t)$ to the CDPM $\mathscr{L}^{\vartheta}({\mathrm{d}}s,{\mathrm{d}}t)$ on $\mathbb{R}^{2}_{\leq}$, by the translation and scaling relations (1.22)-(1.23), it suffices to show the convergence in a restricted domain $[0,1]^{2}_{\leq}:=\{(s,t):0\leq s\leq t\leq 1\}$. For the latter convergence, we prove it in two steps, as detailed below.
In the first step, our objective is to establish Theorem 1.6 below, which ensures that the convergence of random measures defined through the pinning model can be reduced to the convergence of random measures defined through a variant of directed polymer model (for more information about the directed polymer model, see [Com17]). Specifically, we recall the partition function $Z_{M,K}^{\omega,\beta_{N}}$ from (1.14) and the random measure ${\mathcal{Z}}_{N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t)$ from (1.15). We introduce the companion point-to-point directed polymer partition function without boundary disorders
(1.25) $\widetilde{Z}_{M,K}^{\omega,\beta_{N}}(x,y):={\mathbf{E}}\Big{[}\exp\Big{(}\sum\limits_{n=M+1}^{K-1}(\beta_{N}\omega_{n}-\lambda(\beta_{N}))\mathbbm{1}_{\{S_{n}=0\}}\Big{)}\mathbbm{1}_{\{S_{K}=y\}}\Big{|}S_{M}=x\Big{]}$
@@
(1.26) $\widetilde{\mathcal{Z}}_{\lfloor sN\rfloor,\lfloor tN\rfloor}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y):=\sqrt{N}\widetilde{Z}_{\lfloor sN\rfloor,\lfloor tN\rfloor}^{\omega,\beta_{N}}\Big{(}\lfloor x\sqrt{N}\rfloor,\lfloor y\sqrt{N}\rfloor\Big{)}{\mathrm{d}}x{\mathrm{d}}y,$
where the disorder of the directed polymer is only located on the line ×${\mathbbm{Z}}\times\{0\}$. Similar to (1.17), we regard $\widetilde{\mathcal{Z}}_{\lfloor sN\rfloor,\lfloor tN\rfloor}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y)$ as a random variable taking values in the space of locally finite measures on ${\mathbbm{R}}\times{\mathbbm{R}}$ equipped with the topology of vague convergence.
@@
Let $S$ and $\beta_{N}$ be the same as in Theorem 1.1. The following are equivalent.
@@
    The disordered pinning measure ${\mathcal{Z}}_{N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t)$ on $[0,1]^{2}_{\leq}$ defined in (1.15) converges weakly to the unique CDPM $\mathscr{L}^{\vartheta}({\mathrm{d}}s,{\mathrm{d}}t)$ on $[0,1]^{2}_{\leq}$.
@@
    The directed polymer measure $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y)$ on ${\mathbbm{R}}\times{\mathbbm{R}}$ defined in (1.26) converges weakly to a unique limiting random measure ~$\widetilde{\mathscr{L}}^{\vartheta}_{0,1}({\mathrm{d}}x,{\mathrm{d}}y)$ on ${\mathbbm{R}}\times{\mathbbm{R}}$.
Moreover, $\mathscr{L}^{\vartheta}({\mathrm{d}}s,{\mathrm{d}}t)$ and ~$\widetilde{\mathscr{L}}^{\vartheta}_{0,1}({\mathrm{d}}x,{\mathrm{d}}y)$ satisfies the following relationship:
(1.27) ~$\widetilde{\mathscr{L}}^{\vartheta}_{0,1}({\mathrm{d}}x,{\mathrm{d}}y)=\overline{Q}(x,y){\mathrm{d}}x{\mathrm{d}}y+\iint_{0<s<t<1}Q(x,s)\mathscr{L}^{\vartheta}({\mathrm{d}}s,{\mathrm{d}}t)Q(-y,1-t){\mathrm{d}}x{\mathrm{d}}y,$
where $\overline{Q}(x,y)$ is the probability density at time-space $(1,y)$ of a Brownian motion which starts at time-space (0,x) and does not hit 00 in the time interval $(0,1)$, and $Q(x,s)=\frac{|x|}{\sqrt{2\pi}s^{3/2}}e^{-\frac{x^{2}}{2s}}$ is the probability density that a Brownian motion, which starts at time-space $(0,x)$, first hits 00 at time $s$.
With Theorem 1.6 in hand, the second step to conclude Theorem 1.1 is to show the theorem below.
@@
Let $S$ and $\beta_{N}$ be as specified in Theorem 1.1. The random measure $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y)$, defined in (1.26), converges in distribution to a unique limit ~$\widetilde{\mathscr{L}}_{0,1}^{\vartheta}({\mathrm{d}}x,{\mathrm{d}}y)$. The limit is universal, that is, it dose not depend on the law of $\omega$ except for the assumption (1.1).
@@
As shown in [CSZ23, Theorem 1.1], the above convergence can be extended to convergence in finite dimensional distributions of $(\widetilde{\mathcal{Z}}_{\lfloor sN\rfloor,\lfloor tN\rfloor}^{\beta_{N}})_{0\leq s<t<\infty}$ to (~$(\widetilde{\mathscr{L}}^{\vartheta}_{s,t})_{0\leq s<t<\infty}$.
@@
The limiting measures $\mathscr{L}({\mathrm{d}}s,{\mathrm{d}}t)$ and ~$\widetilde{\mathscr{L}}_{s,t}({\mathrm{d}}x,{\mathrm{d}}y)$ corresponds to the point-to-point partition function. The original pinning model (1.12) and directed polymer (1.25), as well as the solutions (1.35) of stochastic Volterra equations and stochastic heat equations to be discussed in the next subsection, are given in terms of point-to-line partition function. It is not hard to see that the scaling limits of these point-to-line (and similarly, line-to-point) partition functions can be obtained by integrating out one relevant coordinate in $\mathscr{L}({\mathrm{d}}s,{\mathrm{d}}t)$ or ~$\widetilde{\mathscr{L}}_{s,t}({\mathrm{d}}x,{\mathrm{d}}y)$.
The advantage of working with the directed polymer measure $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y)$ in Theorem 1.7, rather than the disordered pinning measure ${\mathcal{Z}}_{0,N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t)$ in Theorem 1.1, lies in the introduction of spatial information, resulting in a Markovian structure. Our proof of random measure convergence involves a high moment computation, achieved by coupling multiple polymer chains in the same random environment and then integrating out the environment. Since the random environment is concentrated on the spatial point 0, it is convenient to decompose the polymer chains at each renewal time that anyone of the polymers hits 0. Keeping track of the spatial positions of all polymers, unlike the renewal processes in the pinning model, gives rise to a Markov chain.
1.3. Stochastic Volterra equation and stochastic heat equation
The stochastic Volterra equation has recently attracted vast attention from mathematical finance community. It serves as a rough volatility model, capturing certain key features of short time behavior for asset pricing, see Bayer et. al. [${}^{+}$20] and the references therein. For a smooth function $f$, a parameter $H\in[0,1/2)$, a Brownian motion $W_{\cdot}$ and a random initial condition $\eta$ independent of $W_{\cdot}$, the Volterra equation is defined by
(1.28) $Y_{t}=\eta+\int_{0}^{t}\frac{f(Y_{s})}{(t-s)^{\frac{1}{2}-H}}{\mathrm{d}}W_{s},$
which is deemed rough due to the presence of the singular kernel $(t-s)^{H-\frac{1}{2}}$. In practice, there is evidence [NR18] suggesting that the volatility aligns well with (1.28) for $H$ close to 0.
For constant function $f(x)\equiv c$ and $H>0$, (1.28) is the additive Volterra equation, and its solution $Y^{H}$ is the Riemann-Liouville fractional Brownian motion with Hurst index $H$. Letting $H\downarrow 0$, [NR18] showed that a sequence of suitably rescaled $Y^{H}$ converges to a log-correlated Gaussian measure $Y^{0}$, and thus $Y^{0}$ can be viewed as the solution of additive Volterra equation with zero Hurst index.

### v3

@@ 1.2 The pinning model and weak coupling limit
${\mathbbm{E}}[\zeta_{n}]=0,\quad\sigma_{N}^{2}={\rm Var}(\zeta_{n}):=e^{\lambda(2\beta_{N})-2\lambda(\beta_{N})}-1.$ (1.19)
@@
$\sigma_{N}^{2}=e^{\lambda(2\beta_{N})-2\lambda(\beta_{N})}-1=\frac{1}{R_{N}}\Big(1+\frac{\vartheta+o(1)}{\log N}\Big),\quad\text{for some fixed }\vartheta\in{\mathbbm{R}}.$ (1.20)
Using $\lambda(\beta)=\sum_{i=2}^{4}\frac{\kappa_{i}}{i!}\beta^{i}+O(\beta^{5})$ as $\beta\to 0$, where $\kappa_{i}$ is the $i$-th cumulant of $\omega$, a direct computation (see [13, Appendix A.4] and [15, equation (3.12)]) yields that
$\beta_{N}^{2}=R_{N}^{-1}+C_{1}(\omega)R_{N}^{-3/2}+C_{2}(\omega,\vartheta)R_{N}^{-2}+o(R_{N}^{-2}),$ (1.21)
@@
It is straightforward to see that the CDPM $\mathscr{L}^{\vartheta}$ is translation invariant in law,
$\mathscr{L}^{\vartheta}({\mathrm{d}}(s+{\rm a}),{\mathrm{d}}(t+{\rm a}))\overset{{\rm dist}}{=}\mathscr{L}^{\vartheta}({\mathrm{d}}s,{\mathrm{d}}t),\quad\forall\hskip 1.0pt{\rm a}\in\mathbb{R}.$ (1.22)
@@
$\mathscr{L}^{\vartheta}({\mathrm{d}}({\rm a}s),{\mathrm{d}}({\rm a}t))\overset{{\rm dist}}{=}{\rm a}\hskip 1.0pt\mathscr{L}^{\vartheta+\log{\rm a}}({\mathrm{d}}s,{\mathrm{d}}t),\quad\forall\hskip 1.0pt{\rm a}>0.$ (1.23)
The above translation and scaling relations are analogous to those satisfied by the critical 2d stochastic heat flow [15, Theorem 9.2], and their proofs are exactly the same, which we leave to the readers.
@@
While the first and second moments assumptions on $S$ in (1.10) are standard, the finite $(2+\kappa)$-th moment assumption is for technical reason. In Sections 4 and 5 ((in particular, see (4.26) and (5.13))) we demonstrate that this condition is nearly optimal within our framework. Specifically, to approximate the random walk kernel by the heat kernel, we require more than an $o(1)$ error term in the local limit theorem, namely we need some finer control over the error. Achieving such control necessitates a stronger assumption than the existence of a finite second moment ((see Theorem B.1)).
A general renewal process of the form (1.4) can be generated by Bessel random walks, which are a family of one-dimensional nearest-neighbor random walks with inhomogeneous increments (see [3] for the choice of spatially dependent transition probabilities). For the critical case $\alpha=\frac{1}{2}$ with non-constant slowly varying function $L(\cdot)$, the transition probabilities are slightly inhomogeneous, where
@@
It is also worth noting that, as demonstrated in Tsai [36] very recently for the stochastic heat flow, one may establish a set of conditions that fully characterizes the CDPM. This provides an alternative approach to show the convergence for general underlying renewal processes, which avoids using detailed structures such as the Markov property. It requires only verifying that the discrete model satisfies these conditions in the limit. We plan to pursue this direction in future work.
@@
For two random walks $S$ and $\widetilde{S}$ both satisfying (1.10), it is not hard to see that the difference of their expected number of the overlap $R_{N}-\widetilde{R}_{N}$ has a finite limit as $N$ tends to $\infty$. Denote the limit ${\rm a}({S,\widetilde{S}}):=\lim_{N\to\infty}(R_{N}-\widetilde{R}_{N})$. Suppose $\mathscr{L}^{\vartheta}({\mathrm{d}}s,{\mathrm{d}}t)$ is the limit w.r.t. $S$ and a choice of $\beta_{N}$ satisfying (1.20). Then, for $\widetilde{S}$ and the same $\beta_{N}$,
$\widetilde{\sigma}_{N}^{2}\!=\!\frac{1}{\widetilde{R}_{N}}\Big(1+\frac{\widetilde{R}_{N}-R_{N}}{\log N/(2\pi)}\hskip 1.0pt\frac{\log N/(2\pi)}{R_{N}}\Big)\!\Big(1+\frac{\vartheta+o(1)}{\log N}\Big)\!=\!\frac{1}{\widetilde{R}_{N}}\Big(1+\frac{\vartheta-2\pi{\rm a}({S,\widetilde{S}})+o(1)}{\log N}\Big),$ (1.24)
@@
To show the convergence of ${\mathcal{Z}}_{N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t)$ to the CDPM $\mathscr{L}^{\vartheta}({\mathrm{d}}s,{\mathrm{d}}t)$ on $\mathbb{R}^{2}_{\leq}$, by the translation and scaling relations (1.22)-(1.23), it suffices to show the convergence in a restricted domain $0,1]^{2}_{\leq}:=\{(s,t):0\leq s\leq t\leq 1\}$. For the latter convergence, we prove it in two steps, as detailed below.
In the first step, our goal is to establish Theorem 1.6 below, which ensures that the convergence of random measures defined through the pinning model can be reduced to the convergence of random measures defined through a variant of the directed polymer model (for more information about the directed polymer model, see [17]). Specifically, we recall the partition function $Z_{M,K}^{\omega,\beta_{N}}$ from (1.14) and the random measure ${\mathcal{Z}}_{N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t)$ from (1.15). We introduce the companion point-to-point directed polymer partition function without boundary disorders
$\widetilde{Z}_{M,K}^{\omega,\beta_{N}}(x,y):={\mathbf{E}}\Big[\exp\Big(\sum\limits_{n=M+1}^{K-1}(\beta_{N}\omega_{n}-\lambda(\beta_{N}))\mathbbm{1}_{\{S_{n}=0\}}\Big)\mathbbm{1}_{\{S_{K}=y\}}\Big|S_{M}=x\Big]$ (1.25)
@@
$\widetilde{\mathcal{Z}}_{\lfloor sN\rfloor,\lfloor tN\rfloor}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y):=\sqrt{N}\widetilde{Z}_{\lfloor sN\rfloor,\lfloor tN\rfloor}^{\omega,\beta_{N}}\Big(\lfloor x\sqrt{N}\rfloor,\lfloor y\sqrt{N}\rfloor\Big){\mathrm{d}}x{\mathrm{d}}y,$ (1.26)
where the disorder of the directed polymer is only located on the line ${\mathbbm{Z}}\times\{0\}$. Similar to (1.17), we regard $\widetilde{\mathcal{Z}}_{\lfloor sN\rfloor,\lfloor tN\rfloor}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y)$ as a random variable taking values in the space of locally finite measures on $\mathbbm{R}$$\times{\mathbbm{R}}$ equipped with the topology of vague convergence.
@@
Let $S$ and $\beta_{N}$ be as specified in Theorem 1.1. The following are equivalent.
@@
    The disordered pinning measure ${\mathcal{Z}}_{N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t)$ on $0,1]^{2}_{\leq}$ defined in (1.15) converges weakly to the unique CDPM $\mathscr{L}^{\vartheta}({\mathrm{d}}s,{\mathrm{d}}t)$ on $0,1]^{2}_{\leq}$.
@@
    The directed polymer measure $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y)$ on $\mathbbm{R}$$\times{\mathbbm{R}}$ defined in (1.26) converges weakly to a unique limiting random measure $\widetilde{\mathscr{L}}^{\vartheta}_{0,1}({\mathrm{d}}x,{\mathrm{d}}y)$ on $\mathbbm{R}$$\times{\mathbbm{R}}$.
Moreover, $\mathscr{L}^{\vartheta}({\mathrm{d}}s,{\mathrm{d}}t)$ and $\widetilde{\mathscr{L}}^{\vartheta}_{0,1}({\mathrm{d}}x,{\mathrm{d}}y)$ satisfies the following relationship:
$\widetilde{\mathscr{L}}^{\vartheta}_{0,1}({\mathrm{d}}x,{\mathrm{d}}y)=\overline{Q}(x,y){\mathrm{d}}x{\mathrm{d}}y+\iint_{0<s<t<1}Q(x,s)\mathscr{L}^{\vartheta}({\mathrm{d}}s,{\mathrm{d}}t)Q(-y,1-t){\mathrm{d}}x{\mathrm{d}}y,$ (1.27)
where $\overline{Q}(x,y)$ is the probability density at time-space $(1,y)$ of a Brownian motion which starts at time-space $(0,x)$ and does not hit 0 in the time interval $(0,1)$, and $Q(x,s)=\frac{|x|}{\sqrt{2\pi}s^{3/2}}e^{-\frac{x^{2}}{2s}}$ is the probability density that a Brownian motion, which starts at time-space $(0,x)$, first hits 0 at time $s$.
With Theorem 1.6 in hand, to conclude Theorem 1.1, it suffices to show the theorem below.
@@
Let $S$ and $\beta_{N}$ be as specified in Theorem 1.1. The random measure $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y)$ defined in (1.26) converges in distribution to a unique limit $\widetilde{\mathscr{L}}_{0,1}^{\vartheta}({\mathrm{d}}x,{\mathrm{d}}y)$. The limit is universal, that is, it dose not depend on the law of $\omega$ except for the assumption (1.1).
@@
As shown in [15, Theorem 1.1], the above convergence can be extended to convergence in finite dimensional distributions of $(\widetilde{\mathcal{Z}}_{\lfloor sN\rfloor,\lfloor tN\rfloor}^{\beta_{N}})_{0\leq s<t<\infty}$ to $(\widetilde{\mathscr{L}}^{\vartheta}_{s,t})_{0\leq s<t<\infty}$.
@@
The limiting measures $\mathscr{L}({\mathrm{d}}s,{\mathrm{d}}t)$ and $\widetilde{\mathscr{L}}_{s,t}({\mathrm{d}}x,{\mathrm{d}}y)$ corresponds to the point-to-point partition function. The original pinning model (1.12) and the directed polymer model (1.25), as well as the solutions (1.35) of stochastic Volterra equations and stochastic heat equations to be discussed in the next subsection, are given in terms of point-to-line partition function. It is not hard to see that the scaling limits of these point-to-line (and similarly, line-to-point) partition functions can be obtained by integrating out one relevant coordinate in $\mathscr{L}({\mathrm{d}}s,{\mathrm{d}}t)$ or $\widetilde{\mathscr{L}}_{s,t}({\mathrm{d}}x,{\mathrm{d}}y)$.
The advantage of working with the directed polymer measure $\widetilde{\mathcal{Z}}_{0,N}^{\beta_{N}}({\mathrm{d}}x,{\mathrm{d}}y)$ in Theorem 1.7, rather than the disordered pinning measure ${\mathcal{Z}}_{0,N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t)$ in Theorem 1.1, lies in the introduction of spatial information, resulting in a Markovian structure. Our proof of random measure convergence involves high moments computations, achieved by coupling multiple polymer chains in the same random environment and then integrating out the environment. Since the random environment is concentrated on the spatial point 0, it is convenient to decompose the polymer chains at each renewal time that anyone of the polymers hits 0. Keeping track of the spatial positions of all polymers, unlike the renewal processes in the pinning model, gives rise to a Markov chain.
1.3 Stochastic Volterra equation and stochastic heat equation
The stochastic Volterra equation has recently attracted vast attention from mathematical finance community. It serves as a rough volatility model, capturing certain key features of short time behavior for asset pricing, see Bayer et. al. [5] and the references therein. For a smooth function $f$, a parameter $H\in[0,\frac{1}{2})$, a Brownian motion $W_{\cdot}$ and a random initial condition $\eta$ independent of $W_{\cdot}$, the Volterra equation is defined by
$Y_{t}=\eta+\int_{0}^{t}\frac{f(Y_{s})}{(t-s)^{\frac{1}{2}-H}}{\mathrm{d}}W_{s},$ (1.28)
which is deemed rough due to the presence of the singular kernel $(t-s)^{H-\frac{1}{2}}$. In practice, there is evidence [32] suggesting that the volatility aligns well with (1.28) for $H$ close to 0.
For constant function $f(x)\equiv c$ and $H>0$, (1.28) is the additive Volterra equation, and its solution $Y^{H}$ is the Riemann-Liouville fractional Brownian motion with Hurst index $H$. Letting $H\downarrow 0$, [32] showed that a sequence of suitably rescaled $Y^{H}$ converges to a log-correlated Gaussian measure $Y^{0}$, and thus $Y^{0}$ can be viewed as the solution of the additive Volterra equation with zero Hurst index.